# Language Models can Self-Improve at State-Value Estimation for Better Search

### v2

@@ Acknowledgments
We would like to thank Microsoft’s Azure Accelerate Foundation Models Research Program and NVIDIA’s Academic Grant Program for providing computational resources to support this work. This research is supported in part by the NSF under grant number IIS-2052498 and SMA-2418946. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
Appendix A Self-Taught Lookahead Algorithm
@@
Algorithm 1 Self-Taught Lookahead
0:  Set of tasks $\mathcal{D}_{\texttt{rollout}}$, Base LLM $M$, num iterations $n$, num tasks per iteration $m$
  $\pi_{\theta}\leftarrow\texttt{initialize{\textunderscore}policy{\textunderscore}model}(M)$
  $V_{\phi_{0}}\leftarrow\texttt{initialize{\textunderscore}value{\textunderscore}model}(M)$
  $\mathcal{D}_{\texttt{Train}_{0}}\leftarrow\{\}$
  for $k\leftarrow 1$ to $n$ do
     $\mathcal{D}_{\texttt{rollout}_{k}}\leftarrow\mathcal{D}_{\texttt{rollout}}[m\cdot(k-1):m\cdot k]$Select tasks for iteration $k$}
     for $x\in\mathcal{D}_{\texttt{rollout}_{k}}$ do
        $T\leftarrow\texttt{rollout\_tree}(\pi_{\theta},V_{\phi_{k-1}},x)$Generate rollout search tree for task $x$}
        $y\leftarrow\texttt{calculate\_lookahead\_values}(T)$
        $o\leftarrow\texttt{calculate\_action\_outcomes\_rationales}(T)$
        $y_{\texttt{filtered}},o_{\texttt{filtered}}\leftarrow\texttt{task\_specific\_filter}(y,o)$If applicable, apply task-specific filtering}
        $\mathcal{D}_{\texttt{Train}_{k}}\leftarrow\texttt{add\_new\_data}(\mathcal{D}_{\texttt{Train}_{k}},(y_{\texttt{filtered}},o_{\texttt{filtered}}))$
     end for
     $V_{\phi_{k}}\leftarrow\texttt{fine\textunderscore tune}(V_{\phi_{0}},\mathcal{D}_{\texttt{Train}_{k}})$Finetune from base model}
  end for
Appendix B Broader Impacts
As with all self-improvement work [43, 61], having models self-improve without human supervision may enable agents to learn how to take actions towards task completion that are not well-aligned with human values or preferences. To prevent misuse, we only release code to reproduce experiments in this paper and not to improve general-purpose agents for which these risks might be more prevalent. While investigating these harms is out of the scope of this work, we encourage future research in this area.
Appendix C Limitations
During self-improvement, STL does require task specifications like “I am looking for a queen-sized bed that is black …’ (see Figure 2) for web tasks. However, this assumption is well-founded as prior self-improvement work, such as that on LLM alignment [62, 54], also assumes tasks, or in their case, user prompts, are provided to initialize the self-improvement process. Additionally, as mentioned throughout the paper, this setting is much more reasonable than Reward and Demo Learning and Reward-Guided Inference settings, which require ground truth reward and/or human demonstrations along with task specifications.
@@ D.2 Prompts
Likewise, the prompt used to evaluate states is presented in Figure 9. Note that this evaluation prompt is only used to prompt base models; STL value models are only prompted with the current trajectory. We note that the Likert scale used was crucial to obtaining consistent value outputs on which we could perform STL. For all value estimates (base model or fine-tuned), we prompt the value model 5 times and use the average score as the state value estimate. During the data generation phase, since we need a single rationale to fine-tune on which to construct the action-outcome rationale, we choose the rationale corresponding to the median of the 5 scores.
@@ F.1 Game-of-24 Task
As introduced by [59], the Game-of-24 is a mathematical reasoning task that involves combining four numbers e.g. “2 3 4 5” together with mathematical operations i.e. $+$, --, ///, $\times$ in order to obtain 24. An action in this task consists of simply applying a mathematical operation to combine two numbers e.g. $2+3=5$, the resulting state from the operation is the set of remaining numbers e.g “5 4 5”.
@@ F.4 Implementing STL
Since the state space is quite limited, we combine training examples from the previous $k-1$ iterations with current examples to train the value model in the $k$th iteration. If the same state is encountered multiple times in different iterations, we defer to the value judgment from the latest iteration.
@@ Appendix G Costs
G.1 Inference Costs
To compute inference cost, we use OpenAI’s 444openai.com/api/pricing for OpenAI models like gpt-3.5-turbo and gpt-4o and Groq’s 555groq.com/pricing pricing tables for closed source llama models. Note that we only use Groq for a pricing estimate, as inference and training were run on in-house GPUs. As costs may change over time, we provide the pricing figures that we used for our experiments in Table 4.
@@ G.1 Inference Costs
G.2 Fine-tuning Costs
We also account for the costs incurred from fine-tuning llama-3.1-8b-instruct during STL. Fine-tuning with a single A40 GPU takes 4.5 hours for the WebShop task. While it is difficult to estimate costs on our own in-house GPUs, we can estimate the cost by using VastAI’s figure of \$0.39 per A40 GPU hour  [^6] https://vast.ai/pricing/gpu/A40 when these experiments were run in April, 2025. Thus, the total training run cost roughly \$1.76. We also note that as the number of tasks seen at test time scales, the fine-tuning costs become increasingly negligible.
@@ Appendix H Model Fine-tuning and Serving
Fine-tuning the value model for STL is carried out on a single NVIDIA A40 GPU. We use LoRA finetuning [23] and use models provided by unsloth777unsloth.ai. The hyperparameters used are in Table 5. We fine-tuned Game-of-24 value models for 10 epochs and WebShop value models for 20 epochs due to the differences in difficulty for models to learn the format of the action and state representations.
Additionally, we serve base and fine-tuned models using vLLM 888docs.vllm.ai/en/latest/ [31] for efficient value estimation of new states during search.
Finally, we perform a small analysis on the effect of the discount factor $\gamma$ on search performance. The results of the analysis with a gpt-3.5-turbo policy on WebShop are presented in Table 6. These results confirm the choice of $\gamma=1.0$ from prior work cited in §3.

### v3

@@ Acknowledgments
We would like to thank Microsoft’s Azure Accelerate Foundation Models Research Program and NVIDIA’s Academic Grant Program for providing computational resources to support this work. This research is supported in part by the NSF under grant number IIS-2052498 and SMA-2418946. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. We also appreciate Jungsoo Park, Yao Dou, Geyang Guo, and the anonymous NeurIPS reviewers for their valuable feedback, which helped to improve the paper.
Appendix A Broader Impacts
As with all self-improvement work [45, 66], having models self-improve without human supervision may enable agents to learn how to take actions towards task completion that are not well-aligned with human values or preferences. To prevent misuse, we only release code to reproduce experiments in this paper and not to improve general-purpose agents for which these risks might be more prevalent. While these harms are out of the scope of this work, we encourage future research in this area.
Finally, we note that it is possible to use our value model as a sort of policy to directly select actions in environments with a finite set of actions from each state. While we did not explore how well such a policy would work, improving value estimation to improve the policy may be an interesting direction for future work.
Appendix B Limitations
During self-improvement, STL does require task specifications like “I am looking for a queen-sized bed that is black …” (see Figure 2) for web tasks. However, this assumption is well-founded as prior self-improvement work, such as that on LLM alignment [67, 59], also assumes tasks, or in their case, user prompts, are provided to initialize the self-improvement process. Additionally, as mentioned throughout the paper, this setting is much more reasonable than Reward and Demo Learning and Reward-Guided Inference settings, which require ground truth reward and/or human demonstrations along with task specifications.
STL requires actions to be grounded in language to be able to leverage LLMs’ strong priors. Therefore, STL may struggle in traditional RL tasks like gridworlds, which may have a large state and action space and transitions that are not grounded in language. However, prior work has contextualized gridworld environments by adding landmarks [11], so similar techniques may be used to obtain the necessary consistent and grounded transitions for STL to be successful.
Algorithm 1 Self-Taught Lookahead
1:Set of tasks $\mathcal{D}_{\texttt{rollout}}$, Base LLM $M$, num iterations $n$, num tasks per iteration $m$
[…]
5:for $k\leftarrow 1$ to $n$ do
6:  $\mathcal{D}_{\texttt{rollout}_{k}}\leftarrow\mathcal{D}_{\texttt{rollout}}[m\cdot(k-1):m\cdot k]$ $\triangleright$ Select tasks for iteration $k$
7:  for $x\in\mathcal{D}_{\texttt{rollout}_{k}}$ do
8:   $T\leftarrow\texttt{rollout\_tree}(\pi_{\theta},V_{\phi_{k-1}},x)$ $\triangleright$ Generate rollout search tree for task $x$
9:   $y\leftarrow\texttt{calculate\_lookahead\_values}(T)$
10:   $o\leftarrow\texttt{calculate\_action\_outcomes\_rationales}(T)$
11:   $y_{\texttt{filtered}},o_{\texttt{filtered}}\leftarrow\texttt{task\_specific\_filter}(y,o)$ $\triangleright$ Apply task-specific filtering if applicable
12:   $\mathcal{D}_{\texttt{Train}_{k}}\leftarrow\texttt{add\_new\_data}(\mathcal{D}_{\texttt{Train}_{k}},(y_{\texttt{filtered}},o_{\texttt{filtered}}))$
13:  end for
14:  $V_{\phi_{k}}\leftarrow\texttt{fine\_tune}(V_{\phi_{0}},\mathcal{D}_{\texttt{Train}_{k}})$ $\triangleright$ Finetune from base model
15:end for
Appendix C Algorithms
@@
Additionally, the algorithm for greedy search as used in §4 is presented in Algorithm 2.
@@ D.1 WebShop Task
Algorithm 2 Greedy Search
1:LLM policy $\pi_{\theta}(\cdot)$, LLM value model $V_{\phi}(\cdot)$, Initial state $s_{o}$
[…]
3:while $s_{i}$ is not terminal do
4:  $a_{i}\leftarrow\operatorname*{arg\,max}_{a\in A_{s_{i}}}\Big\{\text{Val}(V_{\phi}(x,s_{0},\ldots,T(s_{i},a))\Big\}$ $\triangleright$ Greedily pick best action
5:  $s_{i+1}\leftarrow T(s_{i},a_{i})$
6:  $i\leftarrow i+1$
7:end while
@@ D.2 Prompts
Likewise, the prompt used to evaluate states is presented in Figure 9. Note that this evaluation prompt is only used to prompt base models; STL value models are only prompted with the current trajectory. We note that the Likert scale used was crucial to obtaining consistent value outputs on which we could perform STL. We also use a special value estimation prompt whenever on a product listing page to select attributes (see Figure 10) in order to obtain consistent values. Specifically, we convert the 4-point scale into a -2 to 2 scale and add the value to the value of the prior state’s (product selection action) value. We tried to use this in our baselines, but we found that it actually harmed the baseline scores. However, in exploratory experiments, it helped yield more consistent values, so we use this for STL.
For all value estimates (base model or fine-tuned), we prompt the value model 5 times and use the average score as the state value estimate. During the data generation phase, since we need a single rationale to fine-tune on which to construct the action-outcome rationale, we choose the rationale corresponding to the median of the 5 scores. After the rebuttal period, we found an instruction to further discount values in the prompts to the fine-tuned models. While we could not rerun all results due to inference costs, we reran the STL evaluation on the mini test set and found results were maintained.
In total, the data generation phase on WebShop yielded a total of 1162 examples, collected from rolling out search trees for 50 tasks. An example of the rationale structure of the training data is presented in Figure 11.
@@
WebShop Attribute Prompt Given an item to purchase and a trajectory that aims to buy an item that exactly matches the specification, analyze how well the last action and observation align with the task. All the last actions you see will be selecting an attribute on the product page of a candidate item. Provide a reflection that concludes with "Thus the correctness score is s", where s is either 1, 2, 3, or 4. Use the following scale for scoring: 1: The attribute selected is opposite to the specified attribute in the task. 2: The attribute selected is irrelevant to the specified attribute in the task. 3: The attribute selected is an attribute mentioned in the instruction, but not all attributes mentioned are currently selected. 4: The attribute selected is an attribute mentioned in the instruction, and all attributes mentioned are currently selected. Keep reflections short (< 100 words). Follow the format of the rationale from the example task below. NOTE: When selecting attributes like click[x-large], click[red], etc., the observations will be “You have clicked x-large” or “You have clicked red” respectively. IMPORTANT: Assume options for size, weight, color, etc. on product pages are not selected unless explicitly mentioned in the trajectory (e.g., click[large], click[red], etc.). Selecting these attributes may be crucial to match the task requirements. Only provide the reflection for the last action. Example Tasks: {few shot examples}  New Task: {input} Respond with the reflection for the last observation of the new task ONLY. As a reminder, the last action and observation is as follows: {last_action} Your response should start with “Reflection:” and end with “Thus the correctness score is …”
Figure 10: Value estimation attribute prompt for WebShop. This prompt was only used to prompt base models during rollouts.
WebShop Rationale Structure Task and Current Trajectory:
Instruction:
i’m looking for a pair of women’s workout shorts with a drawstring waist. i need them to be extra large and in light gray, and price lower than 60.00 dollars [Search]
Action: search[women’s workout shorts drawstring waist extra large light gray < 60 dollars] Observation: removed for brevity Current Action: click[B09T3PJM1R] Current Observation: removed for brevity
STL Output:
I will evaluate the best successor state from the current state:
Best Next Action: click[x02c-gray] Observation of Best Successor State: You have clicked x02c-gray. Reflection of the Best Successor State: The last action selects the color ’x02c-gray’ for the item B09T3PJM1R. Based on the observation, this product’s color is indeed gray, which matches the specified criteria. Therefore, this product matches one of the attributes mentioned in the task. The last action and observation thus capture a step that selects an attribute mentioned in the instruction, but not all attributes mentioned (specifically the size attribute) are currently selected. Thus, the correctness score is 6.00 / 10.00.
Figure 11: Example of rationale structure from the STL training data rolled out on the WebShop task.
@@ D.3 Implementing STL
Finally, we compute loss over both input and output tokens to enable the value model to more quickly capture the dynamics of the environment from the trajectory. However, as we mentioned in §4.1, this practice harms the performance in the setting where the model regresses only on the lookahead value. Therefore, we found that when we only compute loss on the output tokens as in the normal SFT setting, the WebShop score in this setting increases from $70.9$ to $76.0$. This score is actually higher than the setting where we predict the future next best state without a rationale, which indicates that rationales may be key when we try to learn from transition dynamics. We are not sure why this is the case, and why there is not a larger gap between STL and this lookahead value baseline, but we believe that it could be due to small differences in the prompt due to different information settings and/or the prediction or future state potentially harming value estimation for clearly good or bad states.
@@
Table 4: Additional baselines on WebShop. See Table 1 for the full results.
Setting Method Policy Value Mini Test Set (50)
Score $\uparrow$ SR $\uparrow$
ReACT (No Search) gpt-3.5-turbo 68.9 36.0
gpt-3.5-turbo 70.0 36.0
Reward and Demo Free Greedy Baseline gpt-3.5-turbo qwen3-8b 68.6 34.0
Table 5: Score and success rate (SR) on WebShop with error bars. The full results are in Table 1.
Method Policy Value Full Test Set (500)
Score $\uparrow$ SR $\uparrow$
Greedy Baseline gpt-3.5-turbo llama-3.1-8b-instruct 67.7 ± 2.26 26.4 ± 3.86
gpt-3.5-turbo r1-distill-llama-8b 66.3 ± 2.30 24.6 ± 3.78
gpt-3.5-turbo gpt-3.5-turbo 70.6 ± 2.43 35.6 ± 4.23
gpt-3.5-turbo gpt-4o 71.5 ± 2.51 40.6 ± 4.33
gpt-4o llama-3.1-8b-instruct 67.2 ± 2.30 25.8 ± 3.84
gpt-4o r1-distill-llama-8b 66.5 ± 2.33 25.6 ± 3.83
gpt-4o gpt-3.5-turbo 72.4 ± 2.40 38.8 ± 4.29
gpt-4o gpt-4o 71.4 ± 2.49 40.8 ± 4.35
[…]
Human Expert —— —— 82.1 59.6
D.5 Additional Results on WebShop
We have included performance on other baselines in Table 4, including ReACT [65] baselines with OpenAI models and an additional greedy search baseline with a qwen3-8b reasoning value model. Results with error bars are also provided in Table 5.
D.6 Reflexion Performance on WebShop
Tracing bugs in Reflexion implementation
We originally ran Reflexion baselines on WebShop using the official Reflexion GitHub repository [^4] https://github.com/noahshinn/reflexion with three iterations, changing only the model from text-davinci-003 (a deprecated text completion model used in the original paper) to gpt-3.5-turbo. During the author rebuttal period, we found the following:
  • Other researchers had obtained similar Webshop performance to our initial Reflexion results ( 15%) using the unchanged official implementation with gpt-3.5-turbo (see GitHub issue #49).
  • Others also noted that changes to the prompts were needed to adapt the framework to conversational models to see improved performance in the 30-40% range (see GitHub issue #48).
  • Another researcher also identified a bug in the WebShop implementation that prevented the use of memory from prior iterations (see GitHub issue #36). In the discussion of this issue, the first author of the Reflexion paper acknowledged that this bug may have caused the lack of improvement of the Reflexion agent on WebShop that was reported in Appendix B.1 of the original Reflexion paper [49].
After modifying the prompts and patching this memory bug, the ReAct success rate on the 50-task test set is 36%, and the Reflexion success rate is 46% after three iterations, which is similar to the performance reported in other work [73]. We hope that this debugging process can be of some use to the community.
Reflexion vs. LATS.
In Table 1, Reflexion outperforms LATS, which is potentially unexpected as LATS is more computationally expensive than Reflexion. However, due to differences in the mechanisms of these two methods, it is not necessarily the case that LATS performance is lower-bounded by Reflexion performance. For instance, Reflexion enables the LLM policy to make wholesale changes to any steps in its trajectory after reflecting on previous failures. On the other hand, while LATS provides its policy and value LLMs with reflections on previous trajectories, it still relies on the traditional (non-neural) mechanics of MCTS. For example, LLM reflection is not involved during node selection, and instead, the traditional Upper Confidence bounds applied to Trees (UCT) heuristic is used.
@@ E.2 Prompts
In total, the data generation phase on HotpotQA yielded a total of 2708 examples, collected from rolling out search trees for 500 tasks.
@@ F.1 Game-of-24 Task
As introduced by [64], the Game-of-24 is a mathematical reasoning task that involves combining four numbers e.g. “2 3 4 5” together with mathematical operations i.e. ++, -, //, $\times$ in order to obtain 24. An action in this task consists of simply applying a mathematical operation to combine two numbers e.g. $2+3=5$, the resulting state from the operation is the set of remaining numbers e.g “5 4 5”.
@@ F.4 Implementing STL
Since the state space is quite limited, we combine training examples from the previous k1k-1 iterations with current examples to train the value model in the $k$th iteration. If the same state is encountered multiple times in different iterations, we defer to the value judgment from the latest iteration.
@@
Table 6: Performance comparison across different numbers of tasks seen during self-improvement.
Tasks Seen During Improvement Accuracy (25 tasks / iter) Accuracy (50 tasks / iter)
0 38.0 38.0
25 34.0 -
50 30.0 48.0
F.5 Investigating Improvement Dynamics
In §4.3, we claimed that the initial decrease in performance on unseen tasks in Game-of-24 was due to a lack of sufficient tasks seen during self-improvement. We confirmed this claim by rerunning the experiment by rolling out $50$ instead of $25$ tasks per iteration. The results are summarized in Table 6, from which we see that performance actually improves in the first iteration if the number of tasks seen during self-improvement is increased to $50$.
@@ Appendix G Costs
G.1 Data Generation Costs
To compute data generation costs during STL, we use OpenAI’s 555openai.com/api/pricing for gpt-3.5-turbo and Groq’s 666groq.com/pricing pricing tables for closed source llama models. Note that we only use Groq for a pricing estimate, as inference and training were run on in-house GPUs. As costs may change over time, we provide the pricing figures that we used for our experiments in Table 7. We also note that experiments were initially run on gpt-3.5-turbo-0613 on OpenAI Azure, which is multiple times as expensive as the newer gpt-3.5-turbo-0125 due to resource allocation on Azure, despite being the better model. We therefore choose to report the data generation costs using the newer pricing point as it is more comparable to the current prices of other models.
G.2 Fine-tuning Costs
We account for the costs incurred from fine-tuning llama-3.1-8b-instruct during STL. Fine-tuning with a single A40 GPU takes 4.5 hours for the WebShop task. While it is difficult to estimate costs on our own in-house GPUs, we can estimate the cost by using VastAI’s figure of \$0.39 per A40 GPU hour  [^7] https://vast.ai/pricing/gpu/A40 when these experiments were run in April, 2025. Thus, the total training run cost roughly \$1.76. We also note that as the number of tasks seen at test time scales, the fine-tuning costs become increasingly negligible.
G.3 Inference Costs
To compute inference cost, we use the pricing figures in Table 7.
@@ Appendix H Model Fine-tuning and Serving
Fine-tuning the value model for STL is carried out on a single NVIDIA A40 GPU. We use LoRA finetuning [25] and use models provided by unsloth888unsloth.ai. The hyperparameters used are in Table 8. We fine-tuned Game-of-24 value models for 10 epochs and WebShop value models for 20 epochs due to the differences in difficulty for models to learn the format of the action and state representations. We find that we require this large number of epochs to learn both the rationale structure and a good representation of transition dynamics. We note that it is clear that overfitting is now happening since unseen task performance does improve across tasks.
Additionally, we serve base and fine-tuned models using vLLM 999docs.vllm.ai/en/latest/ [33] for efficient value estimation of new states during search. We use $\texttt{temperature}=1.0$ and $\texttt{max\_tokens}=3192$.
As mentioned in §3, we use $\gamma=1.0$ based on prior work. We also ran a small hyperparameter tuning validation for $\gamma\in\{0.95,1.0\}$ on a 50-example held-out validation set and found $\gamma=1.0$ had an average reward of 73.5 compared to 72.7 for $\gamma=0.95$.
Finally, we perform a small analysis on the effect of the discount factor $\gamma$ on search performance on the test. The results of the analysis with a gpt-3.5-turbo policy on WebShop are presented in Table 9. The affect of $\gamma$ on the text set is more stark.